Aluminum has a specific heat of 0.900 J/g·°C. How much energy is required to raise the temperature of 140. g of aluminum from 10.0°C to 232.0°C? Answer in kJ


1
Expert's answer
2021-05-10T01:25:30-0400

Increase in temperature = 232-10 = 222°C

Q = Mc∆T

= 140 × 0.9 × 222

27972J

= 27.972 kJ